Given the system of equations 2 3 2 11 3 2 3 7 4 4 14 x y z x y z x y z + − = − − + = − + = , find the values of x, y, and z using matrix inversion. 


1
Expert's answer
2021-06-07T17:07:00-0400
"2x+3y-2z=11"

"-3x-2y+3z=7"

"4x-4y+z=14"

"A=\begin{pmatrix}
 2 & 3 & -2 \\
 -3& -2 & 3 \\
 4 & -4 & 1\\
\end{pmatrix}"


Augment the matrix with the identity matrix:


"\begin{pmatrix}
 2 & 3 & -2 & & 1 & 0 & 0 \\
 -3 & -2 & 3 & & 0 & 1 & 0 \\
 4 & -4 & 1 & & 0 & 0 & 1\\
\end{pmatrix}"

"R_1=R_1\/2"

"\begin{pmatrix}
 1 & 3/2 & -1 & & 1/2 & 0 & 0 \\
 -3 & -2 & 3 & & 0 & 1 & 0 \\
 4 & -4 & 1 & & 0 & 0 & 1\\
\end{pmatrix}"

"R_2=R_2+3R_1"

"\begin{pmatrix}
 1 & 3/2 & -1 & & 1/2 & 0 & 0 \\
 0 & 5/2 & 0 & & 3/2 & 1 & 0 \\
 4 & -4 & 1 & & 0 & 0 & 1\\
\end{pmatrix}"

"R_3=R_3-4R_1"

"\begin{pmatrix}
 1 & 3/2 & -1 & & 1/2 & 0 & 0 \\
 0 & 5/2 & 0 & & 3/2 & 1 & 0 \\
 0 & -10 & 5 & & -2 & 0 & 1\\
\end{pmatrix}"

"R_2=(2\/5)R_2"

"\begin{pmatrix}
 1 & 3/2 & -1 & & 1/2 & 0 & 0 \\
 0 & 1 & 0 & & 3/5 & 2/5 & 0 \\
 0 & -10 & 5 & & -2 & 0 & 1\\
\end{pmatrix}"

"R_1=R_1-(3\/2)R_2"

"\begin{pmatrix}
 1 & 0 & -1 & & -2/5 & -3/5 & 0 \\
 0 & 1 & 0 & & 3/5 & 2/5 & 0 \\
 0 & -10 & 5 & & -2 & 0 & 1\\
\end{pmatrix}"

"R_3=R_3+10R_2"

"\begin{pmatrix}
 1 & 0 & -1 & & -2/5 & -3/5 & 0 \\
 0 & 1 & 0 & & 3/5 & 2/5 & 0 \\
 0 & 0 & 5 & &4 & 4 & 1\\
\end{pmatrix}"

"R_3=R_3\/5"

"\begin{pmatrix}
 1 & 0 & -1 & & -2/5 & -3/5 & 0 \\
 0 & 1 & 0 & & 3/5 & 2/5 & 0 \\
 0 & 0 & 1 & & 4/5 & 4/5 & 1/5\\
\end{pmatrix}"

"R_1=R_1+R_3"

"\begin{pmatrix}
 1 & 0 & 0 & & 2/5 & 1/5 & 1/5 \\
 0 & 1 & 0 & & 3/5 & 2/5 & 0 \\
 0 & 0 & 1 & & 4/5 & 4/5 & 1/5\\
\end{pmatrix}"


On the left is the identity matrix. On the right is the inverse matrix.



"A^{-1}=\begin{pmatrix}
 2/5 & 1/5 & 1/5 \\
 3/5 & 2/5 & 0 \\
 4/5 & 4/5 & 1/5\\
\end{pmatrix}""AX=B=>A^{-1}AX=A^{-1}B=>X=A^{-1}B"

"X=\begin{pmatrix}
 x\\
 y \\
 z\\
\end{pmatrix}, B=\begin{pmatrix}
 11\\
 7 \\
 14\\
\end{pmatrix}"

"A^{-1}B=\begin{pmatrix}
 2/5 & 1/5 & 1/5 \\
 3/5 & 2/5 & 0 \\
 4/5 & 4/5 & 1/5\\
\end{pmatrix}\begin{pmatrix}
 11\\
 7 \\
 14\\
\end{pmatrix}"

"=\begin{pmatrix}
 (22+7+14)/5\\
 (33+14+0)/5 \\
 (44+28+14)/5\\
\end{pmatrix}=\begin{pmatrix}
 43/5\\
 47/5 \\
 86/5\\
\end{pmatrix}"

"x=\dfrac{43}{5}, y=\dfrac{47}{5}, z=\dfrac{86}{5}"

"(\dfrac{43}{5},\dfrac{47}{5}, \dfrac{86}{5})"